OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable y is eliminated in the first and second equations, then the second and third equations. 3x - y + 3z = 1 4x + y - 6z = 3 4x - 2y + 3z = -3 a) 7x - 3z = 4 12x - 9z = 3 b) 7x - 3z = 4 4x - 6z = 3 c) 3x + 3z = 1 12x - 9z = 3 d) 3x + 3z = 1 4x - 6z = 3

OpenStudy (anonymous):

3x - y + 3z = 1 4x + y - 6z = 3 4x - 2y + 3z = -3 You have these equations and want to eliminate y. By adding the second equation to the first and twice to the third y will be gone. 7x-3z=4 is the first 2 added up

OpenStudy (anonymous):

andras, could help me out real quick?

OpenStudy (anonymous):

now you need to do the second and third (add twice/multiply it by 2 first)

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

What is ti?

OpenStudy (anonymous):

go to the last prob i posted (posted yesterday) and explain to me what i am doing wrong

OpenStudy (anonymous):

OK, one sec

OpenStudy (anonymous):

o k